# Post-detection inference for sequential changepoint localization

### v3

@@ Abstract
This paper addresses a fundamental but largely unexplored challenge in sequential changepoint analysis: conducting inference following a detected change. We construct confidence sets for the unknown changepoint using only the data observed up to a data-dependent stopping time at which an arbitrary sequential detection algorithm $\mathcal{A}$ declares a change. Our framework is nonparametric, makes no assumption on the composite post-change class. We extend it to the composite pre-change class under a suitable assumption. We impose no conditions on the observation space or on $\mathcal{A}$. We thus present the first general tool for sequential changepoint localization, which is theoretically sound and practically effective.
@@ 1 Introduction
$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1.1)
@@ 1.1 Related work
[…]
@@ 1.2 Contributions and paper outline
  • We propose an alternative method of constructing the tests using adaptive thresholds (which is often tighter than the universal threshold), assuming pre- and post-change distributions to be known, in Algorithm 1. We extend this framework to handle parametric pre- and/or post-change classes in Algorithms 2 and 3. These adaptive methods allow the observations to be dependent. We establish corresponding coverage guarantees in Theorems 4.1, 4.4 and 4.5.
@@ 2 Some Fundamental Results
In this section, we highlight some key observations that shape our methodology of post-detection inference for the change-point $T$. We begin by clarifying some notational conventions that will be used throughout the paper.
Notation. For $t\in\mathbb{N}$, let $\mathbb{P}_{F_{0},t,F_{1}}$ denote joint distribution of the sequence $\{X_{n}\}_{n}$ when $X_{1},\dots,X_{t-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$ and $X_{t},X_{t+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{1}$. 333 To understand the key ideas in the paper, it will suffice for the reader to focus on the case when all $X_{i}$’s are all independent, and $F_{0}$ and $F_{1}$ denote the pre- and post-change marginal distributions. However, our framework in Section 4 can actually handle dependent data. Under dependence, $F_{0}$ and $F_{1}$ denotes the joint distributions of $X_{1},\dots,X_{t-1}$ and $X_{t},X_{t+1},\dots$ respectively, and is in fact ambivalent to whether $X_{t}$ is drawn independently of $X_{1},\dots,X_{t-1}$ or conditional on it, so we do not emphasize the distinction in our notation. As a corner cases, let $\mathbb{P}_{F_{0},\infty}$ be the joint distribution when $X_{1},X_{2},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{0}$. In parametric settings, we write $\mathbb{P}_{\theta_{0},t,\theta_{1}}$ and $\mathbb{P}_{\theta_{0},\infty}$ as shorthand for $\mathbb{P}_{F_{\theta_{0}},t,F_{\theta_{1}}}$ and $\mathbb{P}_{F_{\theta_{0}},\infty}$ respectively.
@@
If, for some $\alpha\in(0,1)$, a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}^{*})\geq 1-\alpha$, $\forall T\in\mathbb{N}$, then $\mathbb{P}_{F_{0},\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL is infinite). Said differently, for any algorithm with finite ARL, it is impossible to construct a $1-\alpha$ confidence set that is only a subset of $\{1,\dots,\tau\}$.
Thus, for post-detection inference, the usual notion of coverage is not an appropriate one. Instead, we focus on conditional coverage (1.3), a more reasonable criterion, which was also considered in prior work [5, 36].       The key idea behind constructing our confidence intervals is simple, but several details are subtle. For each candidate $t\in\{1,\cdots,\tau\}$, we test whether $t$ is the true changepoint from some pre-change class $\mathcal{P}_{0}$ to some post-change class $\mathcal{P}_{1}$:
$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\leavevmode\nobreak\ \text{ and }\leavevmode\nobreak\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (2.1)
Let $\phi_{\alpha^{*}}^{(t)}=\mathds{1}(H_{0,t}\text{ is rejected})$ denote some test for the null $H_{0,t}$ such that
$\mathbb{P}_{F_{0},t,F_{1}}(\phi_{\alpha^{*}}^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha^{*},\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$ (2.2)
and form the confidence set by inverting these tests:
@@
[…]
@@
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\leavevmode\nobreak\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.5)
The above result establishes the duality between hypothesis tests and a post-detection confidence set. In the next section, we propose a general framework for constructing confidence sets for $T$ by inverting tests based on the above.
@@ 3 A general framework for nonparametric post-change
$M_{t}=\begin{cases}\prod_{i=t}^{\hat{T}-1}\frac{f_{0}(X_{i})}{f_{1}(X_{i})},&\text{ if }t<\hat{T}\leq\tau<\infty\\ 1,&\text{ if }t=\hat{T}\leq\tau<\infty\\ \prod_{i=\hat{T}}^{t-1}\frac{f_{1}(X_{i})}{f_{0}(X_{i})},&\text{ if }\hat{T}<t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (3.2)
@@
Such powered e-processes for $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ (and vice versa) exist if the total variation distance between the two convex hulls of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ is positive (see e.g., Section 3.4 of [21]). We require all $X_{i}$’s to be independent for both powered (nontrivial) e-processes to be adapted to their respective filtrations.
The final ingredient we need is $\hat{T}$, which is a preliminary point estimate of $T$.  [^4] Here, we again note that our confidence set based on a point estimate $\hat{T}$ can be viewed analogously to the conformal prediction set, which is constructed around a preliminary “point prediction”. For example, one can construct such an estimator based on e-processes under $\mathcal{P}_{0}$. To elaborate, let $\{E^{(t)}_{n}\}_{n\geq t}$ be some forward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$). Then, one reasonable estimate is
$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau}E^{(t)}_{\tau}.$ (3.3)
@@
With the above ingredients in place, we construct a test statistic for $H_{0,t}$ by comparing it against a competing hypothesis $H_{0,T}$. Suppose $t<T$. Then under $H_{0,t}$ (which is false), the segment $X_{t},\dots,X_{T-1}$ follows a post-change distribution from $\mathcal{P}_{1}$ (because under $H_{0,t}$, the change occurred at $t$), whereas under $H_{0,T}$ (which is true), it follows a pre-change distribution from $\mathcal{P}_{0}$. Outside this segment $X_{t},\dots,X_{T-1}$, both hypotheses agree on the distribution of the observations, and hence those are non-informative for testing between $H_{0,t}$ and $H_{0,T}$. Therefore, we use a forward e-process $R^{(t)}_{T-1}$ that accumulates evidence in favor of $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ over this segment. Similarly, if $t>T$, we employ a backward e-process $S^{(t)}_{T}$ to accumulate evidence for $\mathcal{P}_{0}$ against $\mathcal{P}_{1}$. Since the true change-point $T$ is unknown, we replace it with an estimated change-point $\hat{T}$. This naturally leads us to define the following:
$M_{t}=\begin{cases}R^{(t)}_{\hat{T}-1},&\text{ if }t<\hat{T}\leq\tau<\infty\\ 1,&\text{ if }t=\hat{T}\leq\tau<\infty\\ S^{(t)}_{\hat{T}},&\text{ if }\hat{T}<t\leq\tau<\infty\\ -\infty,&\text{ if }t>\tau\text{ or }\tau=\infty.\end{cases}$ (3.4)
@@ 3.1 Confidence set for $T$ with known pre-change distribution
Fix some $N\in\mathbb{N}$. For each $j=1,\dots,N$, simulate a sequence of i.i.d. data points $Z_{1}^{(j)},Z_{2}^{(j)},\dots\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}$ until either the detection algorithm triggers or the sample size reaches $\tau$—whichever occurs first. Let $\tau_{j}:=\min\{\tau,\mathcal{A}(\{Z_{n}^{(j)}\}_{n})\}$ denote the stopping time for the $j$-th simulated sequence. We then estimate $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$ using
$r_{t}=\frac{1}{N}\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t).$ (3.5)
Define the confidence set as
@@
The next theorem proves the validity of the above confidence set.
@@
Suppose that the pre-change distribution is known, i.e., $\mathcal{P}_{0}=F_{0}$. Then, for any $\alpha\in(0,1),T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfying $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (3.6) satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{1}\in\mathcal{P}_{1}.$
Next, we extend it to composite pre-change under an additional assumption.
3.2 A partial extension to composite pre-change

### v4

@@ Abstract
This paper addresses a fundamental but largely unexplored challenge in sequential changepoint analysis: conducting inference following a detected change. We develop a very general framework to construct confidence sets for the unknown changepoint using only the data observed up to a data-dependent stopping time at which an arbitrary sequential detection algorithm declares a change. Our framework is nonparametric, making no assumption on the composite post-change class, the observation space, or the sequential detection procedure used, and is nonasymptotically valid. We also extend it to handle composite pre-change classes under a suitable assumption, and also derive confidence sets for the change magnitude in parametric settings. We provide theoretical guarantees on the width of the confidence intervals. Extensive simulations demonstrate that the produced sets have reasonable size, and slightly conservative coverage. In summary, we present the first general method for sequential changepoint localization, which is theoretically sound and broadly applicable in practice.
@@ 1 Introduction
$X_{1},X_{2},\cdots,X_{T-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle}}{{\sim}}F_{1},$ (1.1)
@@
For instance, rapid shifts in public sentiment often occur on online platforms: a product failure may trigger a sudden surge of negative reviews, or a major event may abruptly alter the tone of social-media discussions. Detecting when such a shift occurs is vital for understanding emerging issues, tracing public reactions to events, or monitoring product reputation in real time. To illustrate the relevance of our methodology in such settings, we conduct an experiment using the Stanford Sentiment Treebank (SST-2) dataset [31], simulating a stream of text reviews whose sentiment abruptly transitions from predominantly positive to predominantly negative. Accurately localizing the changepoint is important: it can help investigate why the change may have occurred.
@@ 1.1 Related work
We are aware of only two directly related works: [5] and [39, Chapter 3]. They also chose to aim for the guarantee (1.3) like us. Both [5, 39] provide asymptotically valid inference after the CUSUM procedure [19] declares a change for known $F_{0}$ and $F_{1}$ in an exponential family; the former only constructs a lower confidence bound for $T$ while the latter also provides a confidence set for $T$. Their approaches have notable methodological and theoretical limitations: (i) they are tailored specifically to the CUSUM procedure, (ii) they assume that the pre- and post-change distributions are both known and they belong to an exponential family, (iii) they only provide asymptotic coverage guarantees (valid when both $T\to\infty$ and ARL $\to\infty$). These constraints limit practical applicability, and their confidence set exhibits undercoverage in our experiments.
@@ 1.2 Contributions and paper outline
  • We propose an alternative method of constructing the tests using adaptive thresholds (which is often tighter than the universal threshold), assuming pre- and post-change distributions to be known, in Algorithm 1. We extend this framework to handle parametric pre- and/or post-change classes in Algorithms 2 and 3. These adaptive methods allow the observations to be dependent. We establish corresponding coverage guarantees in Theorems 4.1, 4.3 and 4.4.
@@ 2 Some Fundamental Results
In this section, we highlight some key observations that shape our methodology of post-detection inference for the change-point $T$.
Notation. For $t\in\mathbb{N}$, let $\mathbb{P}_{F_{0},t,F_{1}}$ denote joint distribution of the sequence $\{X_{n}\}_{n}$ when $X_{1},\dots,X_{t-1}\stackrel{{\scriptstyle}}{{\sim}}F_{0}$ and $X_{t},X_{t+1},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{1}$.  [^3] To understand the key ideas in the paper, it will suffice for the reader to focus on the case when all $X_{i}$’s are all independent, and $F_{0}$ and $F_{1}$ denote the pre- and post-change marginal distributions. However, our framework in Section 4 can actually handle dependent data. Under dependence, $F_{0}$ and $F_{1}$ denotes the joint distributions of $X_{1},\dots,X_{t-1}$ and $X_{t},X_{t+1},\dots$ respectively, and is in fact ambivalent to whether $X_{t}$ is drawn independently of $X_{1},\dots,X_{t-1}$ or conditional on it, so we do not emphasize the distinction in our notation. As a corner cases, let $\mathbb{P}_{F_{0},\infty}$ be the joint distribution when $X_{1},X_{2},\dots\stackrel{{\scriptstyle}}{{\sim}}F_{0}$. In parametric settings, we write $\mathbb{P}_{\theta_{0},t,\theta_{1}}$ and $\mathbb{P}_{\theta_{0},\infty}$ as shorthand for $\mathbb{P}_{F_{\theta_{0}},t,F_{\theta_{1}}}$ and $\mathbb{P}_{F_{\theta_{0}},\infty}$ respectively.
@@
If, for some $\alpha\in(0,1)$, a confidence set $\mathcal{C}^{*}$ that is a subset of $\{1,\cdots,\tau\}$ satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}^{*})\geq 1-\alpha$, $\forall T\in\mathbb{N}$, then $\mathbb{P}_{F_{0},\infty}(\tau=\infty)\geq 1-\alpha$ (meaning that the PFA of the algorithm is at most $\alpha$ and thus the ARL is infinite).
The proposition says that for any algorithm with finite ARL, it is impossible to construct a $1-\alpha$ confidence set that is only a subset of $\{1,\dots,\tau\}$. Thus, for post-detection inference, the usual notion of coverage is not an appropriate one. Instead, we focus on conditional coverage (1.3), a more reasonable criterion, which was also considered in prior work [5, 39]. The key idea behind our methodology is that for each candidate $t\in\{1,\cdots,\tau\}$, we test whether $t$ is the true changepoint from some pre-change class $\mathcal{P}_{0}$ to some post-change class $\mathcal{P}_{1}$:
$H_{0,t}:\;X_{1},\dots,X_{t-1}\sim F_{0}\penalty 10000\ \text{ and }\penalty 10000\ X_{t},X_{t+1},\dots\sim F_{1},\text{ for some }F_{i}\in\mathcal{P}_{i};i=0,1.$ (2.1)
Let $\phi_{\alpha^{*}}^{(t)}=\mathds{1}(H_{0,t}\text{ is rejected})$ denote some test for the null $H_{0,t}$ such that $\mathbb{P}_{F_{0},t,F_{1}}(\phi_{\alpha^{*}}^{(t)}(X_{1},\cdots,X_{\tau})=1)\leq\alpha^{*},\forall F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}$ and form the confidence set by inverting these tests:
@@
In classical statistics, inverting level-$\alpha$ tests yields $1-\alpha$ confidence sets. However, this relationship does not directly hold for us since we test the hypotheses up to a data-dependent time $\tau$, and we are considering conditional coverage.
@@
$\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\penalty 10000\ \forall T\in\mathbb{N},F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}.$ (2.4)
The above result establishes a duality between tests and a post-detection confidence set. Next, we construct the latter by inverting tests based on the above.
@@ 3 A general framework for nonparametric post-change
[…]
@@
Such powered e-processes for $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ (and vice versa) exist if the total variation distance between the two convex hulls of $\mathcal{P}_{0}$ and $\mathcal{P}_{1}$ is positive (see e.g., Section 3.4 of [22]). The likelihood ratio process is the simplest example of an e-process. There is a substantial literature on constructing e-processes in both parametric (e.g., [36]) and nonparametric settings (e.g., [37]). For readers less familiar with these developments, we refer to Chapters 6 and 7 of the recent textbook [22]. Note that we require all $X_{i}$’s to be independent for both powered (nontrivial) e-processes to be adapted to their respective filtrations.
The final ingredient we need is $\hat{T}$, which is a preliminary point estimate of $T$.  [^4] Here, we again note that our confidence set based on a point estimate $\hat{T}$ can be viewed analogously to the conformal prediction set, which is constructed around a preliminary “point prediction”. One can easily construct a point estimate. For example, let $\{S^{(t)}_{n}\}_{n\geq t}$ be some forward $t$-delay e-process under $\mathcal{P}_{0}$ (against $\mathcal{P}_{1}$). Then, one reasonable estimate is
$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau}S^{(t)}_{\tau}$. (3.3)
@@
With the above ingredients in place, we construct a test statistic for $H_{0,t}$ by comparing it against a competing hypothesis $H_{0,T}$. Suppose $t<T$. Then under $H_{0,t}$ (which is false), the segment $X_{t},\dots,X_{T-1}$ follows a post-change distribution from $\mathcal{P}_{1}$ (because under $H_{0,t}$, the change occurred at $t$), whereas under $H_{0,T}$ (which is true), it follows a pre-change distribution from $\mathcal{P}_{0}$. Outside this segment $X_{t},\dots,X_{T-1}$, both hypotheses agree on the distribution of the observations, and hence those are non-informative for testing between $H_{0,t}$ and $H_{0,T}$. Therefore, we use a forward e-process $R^{(t)}_{T-1}$ that accumulates evidence in favor of $\mathcal{P}_{1}$ against $\mathcal{P}_{0}$ over this segment. Similarly, if $t>T$, we use a backward e-process $S^{(t)}_{T}$ to accumulate evidence for $\mathcal{P}_{0}$ against $\mathcal{P}_{1}$, using the segment $X_{T},\dots,X_{t-1}$ . Since the true $T$ is unknown, we replace it with an estimated change-point $\hat{T}$. Fig. 3.1 provides an illustration of the key idea. This naturally leads us to define:
@@
Figure 3.1: Pre-change and post-change parts are shown in black and blue, respectively. Outside the time segment $t,\dots,{\hat{T}-1}]$ when $t<\hat{T}$ (or, $\hat{T},\dots,{t-1}]$ when $t>\hat{T}$) both $H_{0,t}$ and $H_{0,\hat{T}}$ agree on the distribution of the observations.
Although, $M_{t}$ is not an e-variable under $H_{0,t}$ (nor are its constant multiples), thresholding $\frac{M_{t}}{2}$ at $\frac{1}{\alpha}$ yields a level-$\alpha$ test for $H_{0,t}$ — a nontrivial fact shown below.
@@ 3.1 Confidence set for $T$ with known pre-change distribution
Let $r_{t}$ be an unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, that is independent of the data. Then, define the confidence set as
@@
One natural approach to estimate $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$ is via Monte Carlo simulation. Fix $N\in\mathbb{N}$ and for each $j=1,\dots,N$, simulate a sequence of i.i.d. data points $Z_{1}^{(j)},Z_{2}^{(j)},\dots\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}$ until either the detection algorithm triggers or the sample size reaches $\tau$—whichever occurs first. Let $\tau_{j}:=\min\{\tau,\mathcal{A}(\{Z_{n}^{(j)}\}_{n})\}$ denote the stopping time for the $j$-th simulated sequence. Then $r_{t}=\frac{1}{N}\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)$ is a natural unbiased estimator of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$ ,but it can equal zero with positive probability. To avoid this, we fix some $r\geq 2$ and for each $t$, one might define $N_{t}$ to be a random index until $\tau_{j}\geq t$ happens exactly $r$ many times. Then $r_{t}=\frac{r-1}{N_{t}-1}$ is an unbiased estimator, which is strictly positive. Alternatively, one may use an asymptotically unbiased estimator with $N$, such as, $r_{t}=\frac{1}{N+1}\left(1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)\right)$, which guarantees strict positivity but only provides asymptotic $1-\alpha$ coverage guarantee as $N\to\infty$. The next theorem proves the validity of the above confidence set.
@@
Suppose that the pre-change distribution is known, i.e., $\mathcal{P}_{0}=F_{0}$. Then, for any $\alpha\in(0,1),T,N\in\mathbb{N}$, post-change class $\mathcal{P}_{1}$, and change detection method $\mathcal{A}$ with stopping time $\tau$ satisfying $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}$ defined in (3.5), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha,\forall F_{1}\in\mathcal{P}_{1}.$
Now we analyze the size of the confidence set (3.5) for known pre-change $F_{0}$ and post-change $F_{1}$, with $r_{t}=\frac{1+\sum_{j=1}^{N}\mathds{1}(\tau_{j}\geq t)}{N+1}$, $M_{t}$ as defined in (3.2) with $\hat{T}$ as defined in (3.1). We provide a bound on the expected size of the confidence set as $N\to\infty$, where $N$ is the number of simulations used to calculate $r_{t}$. Define, $p_{t}=\mathbb{P}_{F_{0}^{*},\infty}[\tau\geq t]$. Note that $r_{t}\to p_{t}$ as $N\to\infty$. Define $\rho_{0}(s)=\mathbb{E}_{F_{0}}\left(\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\right)^{s}$ and $\rho_{1}(s)=\mathbb{E}_{F_{1}}\left(\frac{f_{0}(X_{i})}{f_{1}(X_{i})}\right)^{s}$. These two quantities measure closeness between $F_{0}$ and $F_{1}$, and thus the problem hardness. Note that for $s=1/2,$ this coincides with Bhattacharyya distance. At $s=0,1$, $\rho_{0}$ and $\rho_{1}$ take value $1.$ Since for $s\in(0,1)$, $h(x)=x^{s}$ is a strictly concave function and since $F_{1}\neq F_{0}$, $\frac{f_{1}(X)}{f_{0}(X)}$ is not almost everywhere constant, we get a strict inequality using Jensen’s: $\rho_{0}(s)=\mathbb{E}_{F_{0}}\left[\left(\frac{f_{1}(X)}{f_{0}(X)}\right)^{s}\right]<\left(\mathbb{E}_{F_{0}}\left[\frac{f_{1}(X)}{f_{0}(X)}\right]\right)^{s}=1.$ Similarly, for $s\in(0,1),\rho_{1}(s)<1$. We now define $s_{0}=\arg\min_{s\in[0,1]}\rho_{0}(s)$ and $s_{1}=\arg\min_{s\in[0,1]}\rho_{1}(s)$, and note that the preceding facts imply that $0<s_{0}<1$ and $0<s_{1}<1$. Our length bound will depend on the expected detection delay, $\Delta_{F_{0},T,F_{1}}:=\mathbb{E}_{F_{0},T,F_{1}}(|\tau-T|\mid\tau\geq T).$ For typical detection algorithms, this quantity often scales like the logarithm of the average run length divided by the KL divergence between $F_{1},F_{0}$. Finally, we call the stopping time $\tau$ as sensitive if
$\penalty 10000\ \mathbb{P}_{F_{0},T,F_{1}}[\tau\leq t]\geq\mathbb{P}_{F_{0},\infty}[\tau\leq t],$ (3.6)
for all $t$ and $T$. Equivalently, $\tau$ under a change is stochastically smaller than $\tau$ under no change. We can expect all practically used detection algorithms to be sensitive, as they should stop later when a change is absent than when it is present.
Theorem 3.3.
In the setting described above, for any change detection method with stopping time $\tau$ satisfying and $p_{T}\neq 0$, the size of $\mathcal{C}$} can be bounded as:
$\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\cdot\frac{\rho_{0}(s_{0})-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{F_{0},T,F_{1}}.$ (3.7)
Additionally, if $\tau$ is sensitive in the sense of (3.6), the last term in (3.7) can be improved to $\Psi_{F_{0},T,F_{1}}:=\min\left\{\Delta_{F_{0},T,F_{1}},\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right)\right\}$.
Remark 3.4.
Suppose $F_{1}=N(\mu_{1},\sigma^{2})$ vs $F_{0}=N(\mu_{0},\sigma^{2})$ and we are using a CUSUM detector with the likelihood ratio (which satisfies the assumption) with average run length $A$. Consider the asymptotic regime when $A\to\infty$, and $T=CA$, for some $C>0$. Using the property of the CUSUM detector that $\frac{\tau}{\mathbb{E}_{F_{0},\infty}(\tau)}\stackrel{{\scriptstyle d}}{{\to}}\exp(1)$ as $A\to\infty$, we have $p_{T}=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)=\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq CA_{T})\to\exp(-C)$. Also, for the Gaussian distribution, it is easy to verify that $\rho_{0}(s)=\rho_{1}(s)=\exp\left(-s(1-s)K\right)$, where $K=\frac{(\mu_{1}-\mu_{0})^{2}}{2\sigma^{2}}$ and $s_{0}=s_{1}=1/2$. Substituting, we get
$\limsup_{T,A\to\infty,T=CA}\limsup_{N\to\infty}\mathbb{E}_{F_{0},T,F_{1}}\left(|\mathcal{C}|\mid\tau\geq T\right)\leq$ $\sqrt{\frac{2}{\alpha}}\left[\frac{\exp(\frac{3C}{2}-\frac{K}{4})}{1-\exp\left(-\frac{K}{4}\right)}+\frac{\exp\left(C-\frac{K}{8}\right)}{1-\exp\left(-\frac{K}{8}\right)}\right]$
$+1+\frac{\exp(C-\frac{K}{4})}{1-\exp\left(-\frac{K}{4}\right)}.$
To simplify further, consider the difficult regime of small $K$, where $\left(\exp\left(\frac{K}{4}\right)-1\right)\approx K/4$, and $\left(\exp\left(\frac{K}{8}\right)-1\right)\approx K/8$, and take $C=\frac{1}{3}\log(\frac{2}{\alpha})$, so that the above expression is bounded by, ignoring constants, $K^{-1}[\frac{1}{\alpha}+\frac{1}{\alpha^{5/6}}++\frac{1}{\alpha^{1/3}}]$. The fact that the bound grows with $C=T/A$ is empirically verified in Supplementary Section G.1. This behavior is not a limitation of our methodology; rather, it reflects a fundamental difficulty of the problem: $p_{t}$ decays as $C$ increases, and hence, by our duality result (Proposition 2.2), the tests corresponding to any confidence set will have smaller levels $\alpha p_{t}$, leading to reduced power and, consequently, wider confidence sets.
3.2 An extension to composite pre-change